OpenStudy (anonymous):

Multiply and simplify if possible. √7x(√x-7√7)

OpenStudy (fools101):

Hi! http://www.mathway.com/problem.aspx?p=algebra let me know i this helps! ~Jess~

OpenStudy (zehanz):

√(7x)(√x-7√7) is what you mean, I think... Use the distributive property: a(b-c)=ab-ac

OpenStudy (anonymous):

no the way I wrote it is how it is on my homework..

OpenStudy (zehanz):

It is also good to know that √a√b=√(ab)

OpenStudy (zehanz):

OK, assuming it is √7*x(√x-7√7), use a(b-c)=ab-ac

OpenStudy (anonymous):

can you show me how?

OpenStudy (anonymous):

okay thank you!

OpenStudy (fools101):

Your welcome ^_^